# Forecasting Arctic Temperatures with Temporally Dependent Data Using Quantile Gradient Boosting and Adaptive Conformal Prediction Regions

### v2

@@ 1 Introduction
There have been recent technical advances (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, equivocal reliability, and other issues remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of legitimate concerns about a warming climate.
@@ 2 Svalbard, Norway: The Forecasting Setting
Snow falls during much of the year, but total precipitation is relatively low – technically making Svalbard a polar desert. From late October to mid-February the archipelago experiences the polar night, when the sun does not rise above the horizon. From mid-April to late August, the sun remains above the horizon, bringing continuous daylight. In recent decades Svalbard has warmed at a rate several times the global average, with mean temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s regions most affected by climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
@@ 2.1 Some Implications for Forecasting Temperatures
A practical attraction of Svalbard for temperature forecasting is the weather station at Longyearbyen Airport, whose data are curated by NOAA in collaboration with its international partners. These data are freely available and easy to download. Under the Svalbard Treaty of 1920, Norway retains sovereignty but grants citizens of other signatory nations equal rights to engage in scientific and commercial activities. International research stations are operated by Russia, Poland, Germany, China, the United States, and others.
@@ 3 Data
Anticipating the use of adaptive conformal prediction regions (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 provide an “honest” assessment of the number of algorithm iterations required. They are also the observations needed to construct proper nonconformal scores.
@@
Predictors are all lagged by 14 days because that seems to be a lag commonly used by climate scientists (Li et al., 2024). In future work, longer lags could be considered. The predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^4] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are therefore included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets chosen at random from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement for the 2022 and 2024 holdout samples is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. After a search in Google Scholar, their combination appear to be novel. Each component is discussed in a grounded manner when deployed. Appendix A provides a step-by-step synopsis.
@@ 5 Results
The results begin with simple univariate plots to provide context for the more involved analyses. Because of the months-long alternation between daylight and darkness, Arctic temperatures fluctuate differently from those at lower latitudes. In general, the figures that follow should be largely self-explanatory.
@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 represents realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations caused by Arctic amplification.
@@ 6 Forecasts and Estimated Uncertainty
Figure 7 illustrates how conformal prediction regions work. The figure highlights important features of the Svalbard data but is not meant to be a reproduction of those data. On the horizontal axis is the forecasted temperature. On the vertical axis is the observed future temperature that would, in practice, be unknown when a forecast is made. The solid blue line conveys how the forecasts and future temperatures are systematically related. The dashed red lines are the upper and lower bounds of the adaptive prediction region for each forecast. If $1-\alpha=0.80$, the prediction region for any forecast contains the correct future temperature with a probability of at least 0.80.
Figure 7: Simulation of adaptive conformal prediction regions from quantile gradient boosting with $\tau=0.80$. The horizontal axis is the forecasted temperature. The vertical axis is the observed future temperature. The solid blue line shows how forecasts and true future temperatures are systematically related. The dashed red lines show the upper and lower bounds of the adaptive prediction region, and the two vertical dotted magenta lines mark the prediction regions when the forecasted temperature happens to be $10^{\circ}\mathrm{C}$ or $5^{\circ}\mathrm{C}$.
To help fix these ideas, consider two illustrative forecasts, one at $10^{\circ}\mathrm{C}$ and the other at $5^{\circ}\mathrm{C}$. Several features of their prediction regions are apparent and apply to all forecasted temperatures in the figure. Forecasts are not necessarily in the middle of their prediction regions, which means that the regions need not be symmetric around the forecast. In addition, because for the quantile gradient-boosting algorithm $\tau$ was set to 0.60, the forecasts tend to fall above each region’s median. Finally, the lengths of the adaptive prediction regions grow shorter as the fit improves (Romano, Patterson and Candès, 2019). In Figure 7, forecasts above $0^{\circ}\mathrm{C}$ fit the observed temperatures better and have shorter region lengths. In the Svalbard setting, forecasts would be produced one day at a time, but Figure 7 is meant to convey the results after many forecasts are made. The figure also makes plain that there is no requirement that the forecasts be correct.
@@ 6.1 Uncertainty Results
The lengths of the 2024 conformal prediction regions vary substantially, consistent with Figure 3. For days in which the forecasted 2 p.m. temperature exceeds 0, the average half-width is about $\pm 2.6^{\circ}\mathrm{C}$, and the half-widths range from about $1.4^{\circ}\mathrm{C}$ to $6.2^{\circ}\mathrm{C}$. For days in which the forecasted 2 p.m. temperature does not exceed 0, the average half-width is about $\pm 4.7^{\circ}\mathrm{C}$, and the half-widths range from about $2.0^{\circ}\mathrm{C}$ to $8.3^{\circ}\mathrm{C}$. Performance is better for the warmer diurnal temperatures because $\tau>0.50$.
@@
Figure 8: For the 2024 holdout sample, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
@@
But physical interpretations can be subtle. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$. Before that stage, a two-phase mixture of nearly pure ice and brine remains. Melting reverses this process: pure ice begins melting at $0^{\circ}\mathrm{C}$, whereas frozen brine melts at lower temperatures.
The rates of melting and freezing depend on cloud cover, winds, and ocean currents that transport warmer or cooler water to local sites. Consequently, freezing and melting occur over a range of temperatures and time scales rather than at a fixed threshold. This interpretation aligns with Figure 8.
@@
Given these results, the role of conformal prediction regions can be briefly revisited. For a new, unlabeled case T+1T{+}1, the true temperature will fall within its adaptive conformal prediction region with probability at least $1-\alpha$. When the lower bound of a prediction region is greater than zero, the probability that the future temperature will be above 0 is $1-\alpha/2$. For a specified coverage probability of 0.80, that probability would increase to 0.90. Coverage is gained because under some circumstances, the prediction region’s upper bound may not matter much. This approach could work well for Svalbard insofar as temperatures $0^{\circ}\mathrm{C}$ are far more concerning than lower temperatures. In the spirit of Figure 8, when the lower bound of the prediction region is greater than $0^{\circ}\mathrm{C}$, one may simply assume that the forecast of melting is correct.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, which is robust to sparse data. To the best of the author’s knowledge, the combined use of three temporally separated samples, quantile gradient boosting, AR(1)-based whitening, and regime-specific adaptive conformal prediction has not previously appeared in the literature. The statistical procedures were informed by extensive knowledge of the research site. Equally important was guidance from climate science in the choice of algorithms and in interpreting their output. The aim was to contribute both to scientific understanding and to practical human adaptations to Arctic melting.
@@
More broadly, forecasts on a two-week timescale could foster coordination among local administrators, regional governments, and research stations, improving situational awareness and resource allocation across sectors. Even modest improvements in lead time may yield substantial adaptive benefits in Arctic regions where melt conditions evolve rapidly and logistical flexibility is limited. [^14] The last three paragraphs were informed by a literature search undertaken by ChatGPT, which also provided initial wording later edited for consistency and style.
@@ 8 Conclusions
Projections of future global warming have long been central to IPCC assessments, complemented by many smaller-scale forecasting efforts. The analyses presented here contribute to the latter. Temperature forecasts with a two-week lead time appear promising, especially when focused on Arctic thresholds associated with widespread melting. Forecasting uncertainty is explicitly addressed, and results are presented in a manner intended to be accessible and useful to stakeholders.
The analyses also illustrate that one need not rely on “industrial-strength” data or computational infrastructure to obtain meaningful forecasts. With thoughtfully applied algorithms and accessible data, informative short-term forecasts can be generated on an ordinary desktop computer with an internet connection.

### v3

@@ 1 Introduction
There have been recent technical advances (Pathak et al., 2022; Price et al., 2024; Bodnar et al., 2025). Yet issues of coarse spatial scales, standardized output, and equivocal reliability remain largely unresolved (Ortega et al., 2022; Balaji et al., 2022; Chang et al., 2023; Zhang et al., 2025). This paper uses Svalbard, Norway, as a test case to advance Arctic temperature forecasting while addressing these three limitations. Temperatures are the focus because of legitimate concerns about a warming climate.
@@ 2 Svalbard, Norway: The Forecasting Setting
Svalbard has warmed at a rate several times the global average, with mean temperatures increasing by more than $5^{\circ}\mathrm{C}$ since the 1970s. Glaciers are retreating, sea ice is thinning, and thawing permafrost is reshaping the terrain. The archipelago is now among the world’s regions most affected by climate change (Urbański and Litwicka, 2022; Karlsen et al., 2024; Bradley et al., 2025; Schuler et al., 2025).
@@ 3 Data
Anticipating the use of adaptive conformal prediction regions (Romano, Patterson and Candès, 2019), a variant on split-sample methods is employed. Observations for 2023 constitute the training data. Calibration data from 2022 can provide, in principle, an “honest” assessment of the number of algorithm iterations required (Hyndman and Athanasopoulos, 2021, sec. 5.8). They are also the observations needed to construct proper nonconformal scores.
@@
Predictors are all lagged by 14 days because that seems to be a lag commonly used by climate scientists (Li et al., 2024). In future work, longer lags could be considered. The weather station predictors include (1) wind direction in degrees from true north, (2) wind speed in meters per second, (3) air temperature in degrees Celsius, (4) atmospheric pressure in hectopascals (hPa), (5) visibility in meters, (6) dew point in degrees Celsius, (7) relative humidity in percent, and (8) a day counter ranging from 1 to 365. The counter can capture temporal trends: on average, the diurnal months are warmer than the nocturnal months, although temperature changes over time may be nonlinear. Several predictors are likely to interact in complex ways (Semenov, 2021). The 14-day lags translate into fitted 2 p.m. temperatures two weeks later that are a foundation for forecasts 14 days in advance. [^4] Lagging variables is a routine procedure in feature construction. In this case, however, lagging the first 14 days of a year pushes their reference points into the final 14 days of the previous year. Data from 2021 are included to provide lagged values for the first 14 days of 2022. Otherwise, the 2021 data are not used.
The complete dataset forms a multiple time series, making temporal dependence a potential complication. Split samples are commonly disjoint subsets chosen at random from the data available, but that is likely to obscure any temporal dependence (Hyndman and Athanasopoulos, 2021, sec. 5.8). A key requirement for the 2022 and 2024 holdout samples is that the same physical processes apply during the identical months of 2022, 2023, and 2024; data for all three years are treated as random realizations from the same underlying joint probability distribution. Nonetheless, the pace of Arctic amplification may affect this comparability, an issue examined empirically in Section 5.2.
@@ 4 Statistical Methods
The analyses to follow produce legitimate forecasts from supervised machine learning applied to temporal data. Holdout samples are used to construct provably valid prediction regions. The statistical framework is an interlocking set of procedures. After a search in Google Scholar, their combination appears to be novel. Each component is discussed in a grounded manner when deployed. Appendix A provides a step-by-step synopsis.
@@ 5 Results
The results begin with simple univariate plots to provide context for the more involved analyses. Because of the months-long alternation between daylight and darkness, Arctic temperatures fluctuate differently from those at lower latitudes. The figures that follow should be largely self-explanatory.
@@ 5.2 Time-Series Plot for Svalbard Temperatures in 2022, 2023, and 2024
Recall the premise that each time series of daily 2 p.m. temperatures for 2022, 2023, and 2024 represents realizations from the same underlying joint probability distribution; the underlying physics for corresponding months in those three years should be comparable. However, there could be aberrations caused by rapid Arctic amplification.
@@ 6 Forecasts and Estimated Uncertainty
Figure 7 illustrates how adaptive conformal prediction regions work. The figure highlights important features of the Svalbard data but is not meant to be a reproduction of those data. It is a didactic device only.
On the horizontal axis is the forecasted temperature. On the vertical axis is the observed future temperature that would, in practice, be unknown when a forecast is made. The solid blue line conveys how the forecasts and future temperatures are systematically related. The dashed red lines are the upper and lower bounds of the adaptive prediction region for each forecast. If $1-\alpha=0.80$, the prediction region for any forecast contains the correct future temperature with a probability of at least 0.80.
Figure 7: A didactive simulation of adaptive conformal prediction regions from quantile gradient boosting with $\tau=0.80$. The horizontal axis is the forecasted temperature. The vertical axis is the observed future temperature. The solid blue line shows how forecasts and true future temperatures are systematically related. The dashed red lines show the upper and lower bounds of the adaptive prediction region, and the two vertical dotted magenta lines mark the prediction regions when the forecasted temperature happens to be $10^{\circ}\mathrm{C}$ or $5^{\circ}\mathrm{C}$.
To help fix these ideas, consider two illustrative forecasts, one at $10^{\circ}\mathrm{C}$ and the other at $5^{\circ}\mathrm{C}$. Several features of their prediction regions are apparent and apply to all forecasted temperatures in the figure. Forecasts are not necessarily in the middle of their prediction regions, which means that the regions need not be symmetric around their forecasts. In addition, because for the quantile gradient-boosting algorithm $\tau$ was set to 0.60, the forecasts tend to fall above each region’s median. Finally, the lengths of the adaptive prediction regions grow shorter as the fit improves (Romano, Patterson and Candès, 2019). In Figure 7, forecasts above $0^{\circ}\mathrm{C}$ fit the observed temperatures better and have shorter region lengths. In the Svalbard setting, forecasts would be produced one day at a time, but Figure 7 is meant to convey the results after many forecasts are made. The figure also makes plain that there is no requirement that the forecasts be correct.
@@ 6.1 Uncertainty Results
For the 2024 test data, the lengths of the conformal prediction regions vary substantially, consistent with Figure 3. For days in which the forecasted 2 p.m. temperature exceeds 0, the average half-width is about $\pm 2.6^{\circ}\mathrm{C}$, and the half-widths range from about $1.4^{\circ}\mathrm{C}$ to $6.2^{\circ}\mathrm{C}$. For days in which the forecasted 2 p.m. temperature does not exceed 0, the average half-width is about $\pm 4.7^{\circ}\mathrm{C}$, and the half-widths range from about $2.0^{\circ}\mathrm{C}$ to $8.3^{\circ}\mathrm{C}$. Performance is better for the warmer diurnal temperatures because $\tau>0.50$.
@@
Figure 8: For the 2024 holdout test sample, the mean forecasted temperature in a bin is on the horizontal axis. The percentage of cases in each of 20 equal-width bins having a true future temperature exceeding $0^{\circ}\mathrm{C}$ is on the vertical axis. Filled circles correspond to bins with mean forecasted temperature above $0^{\circ}\mathrm{C}$. Open circles correspond to those below. The solid blue line is a loess smoother.
@@
But physical interpretations can be subtle, especially for sea ice. When sea ice forms, its hexagonal lattice cannot readily incorporate salt ions. The excluded salt forms brine pockets and channels among the ice grains, leaving the ice as nearly fresh water while the brine becomes increasingly salty. As freezing continues, the concentrated brine’s freezing point can fall well below $0^{\circ}\mathrm{C}$ because the salt ions attract water molecules the lattice needs; it is more difficult for the water molecules to self-organize into the lattice structure. Unlike fresh water ice, which melts at a single fixed temperature, sea ice melts over a range of temperatures, because the brine pockets within it are already liquid at temperatures below $0^{\circ}\mathrm{C}$. As the temperature rises, progressively more of the remaining ice melts, and the brine channels widen and connect. Because a significant fraction of sea ice consists of these pre-existing liquid brine inclusions, it requires less energy to melt than an equivalent mass of fresh water ice — the brine pockets need no latent heat to liquify, only the nearly pure ice fraction does. By the same logic, the energy released when salt water freezes is less than that released when fresh water freezes. The amount of energy required to melt sea ice equals the amount released when it freezes, consistent with thermodynamic reversibility. This interpretation aligns with Figure 8.
@@
Given these results, the role of adaptive conformal prediction regions can be briefly revisited. For a new, unlabeled case T+1T{+}1, the true temperature will fall within its adaptive conformal prediction region with probability at least $1-\alpha$. When the lower bound of a prediction region is greater than zero, the probability that the future temperature will be above 0 is at least $1-\alpha/2$. For a specified coverage probability of 0.80, that probability would be at least 0.90. Coverage is gained because under some circumstances, the prediction region’s upper bound may not matter much. This approach could work well for Svalbard insofar as temperatures $0^{\circ}\mathrm{C}$ are far more concerning than lower temperatures. In the spirit of Figure 8, when the lower bound of the prediction region is greater than $0^{\circ}\mathrm{C}$, one may simply assume that the forecast of melting is correct.
@@ 7 Discussion
The analysis of 2 p.m. temperatures in Svalbard, Norway, combines three interlocking statistical traditions: gradient boosting with a quantile as the estimation target, an AR(1) model to construct exchangeable nonconformal scores, and adaptive conformal prediction regions computed with quantile random forests, which is robust to sparse data. To the best of the author’s knowledge, these procedures applied to three temporally separated samples has not previously appeared in the literature.
@@
More broadly, forecasts on a two-week timescale could foster coordination among local administrators, regional governments, and research stations, improving situational awareness and resource allocation across sectors. Even modest improvements in lead time may yield substantial adaptive benefits in Arctic regions where melt conditions evolve rapidly and logistical flexibility is limited. [^14] The last three paragraphs were informed by a literature search undertaken by ChatGPT, which also provided some initial wording later edited for consistency and style.
@@ 8 Conclusions
Projections of future global warming have long been central to IPCC assessments, complemented by many smaller-scale forecasting efforts. The analyses presented here contribute to the latter. Temperature forecasts with a two-week lead time appear promising, especially when focused on Arctic thresholds associated with widespread melting. Forecasting uncertainty is explicitly addressed, and results are presented in a manner intended to be accessible and useful to stakeholders. With thoughtfully applied algorithms and accessible data, informative short-term forecasts can be generated on an ordinary desktop computer with an internet connection.